OpenStudy (anonymous):

A rectangular garden has a width of 5 feet and a length of 10 feet. If an equal amount is added to both the width and the length, the area is increased to 84 square feet. What is this amount

OpenStudy (anonymous):

84 = 7 * 12, so we can get an area of 84 square feet by adding two feet to the width and the length.

OpenStudy (anonymous):

Let, the equal amout is x Therefore, 2{(5+x)+(10+x)}=84 2(15+2x)=84 4x=84-30 x=54/4 x=13.5

OpenStudy (anonymous):

\[(x+10) \{x+5\}=84 \]x = 2